Sodium cyanide is the salt of the weak acid HCN. Calculate the concentration of H3O^+ , OH^- , HCN and Na^+ in solution prepared by dissolving 10.8 g of NaCN in enough water to make 5.0 x 10^2 mL of solution at 25 degrees Celsius.
1
Expert's answer
2020-11-12T06:38:27-0500

Solution:

Ka of HCN = 4.0×10-10

M of NaCN = 10.8g NaCN/49.01g/mol NaCN/0.5L = 0.441 M

CN - + H2O ↔ HCN + OH-

0.441-x x x

Kb = 1×10-14 / 4.0×10-10 = 2.5×10-5

Kb = [HCN] [OH-] / [CN-]

2.5×10-5 = x2/(0.441-x) ≈ x2/0.441

x2=1.10×10-5

x= 0.0033M

[HCN] = 0.0033M

[OH-] = 0.0033M

[Na+] = 0.44M

[H3O+] = 3.0×10-12M

Answer: [HCN] = 0.0033M; [OH-] = 0.0033M; [Na+] = 0.44M; [H3O+] = 3.0×10-12M